Obtain an expression of probability for finding the free particle at position x
confined in a box of length a.
1
Expert's answer
2020-04-07T09:19:58-0400

The normalized wavefunction for a free particle found at "x" in a 1-dimensional box of length "a" can be found from Schrödinger equation:


"\psi=\sqrt{\frac{2}{a}}\text{sin}\frac{n\pi}{a}x."